OpenStudy (anonymous):

If X = 5 inches and Y = 12 inches, what is the measure of Z? A. 7 inches B. 9 inches C. 11 inches D. 13 inches

OpenStudy (anonymous):

picture below

OpenStudy (cwrw238):

this the pythagoras theorem z^2 = 5^2 + 12^2

OpenStudy (cwrw238):

so z^2 = 25 + 144 = 169 so z = square root of 169 can you finish this off?

OpenStudy (anonymous):

yess I Got it thankss!!

OpenStudy (cwrw238):

yw